OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). x+3y=2 2x+3y=22 Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). x+3y=2 2x+3y=22 @Mathematics

OpenStudy (lalaly):

x=2-3y 2(2-3y)+3y=22 4-6y+3y=22 -3y=18 y=-6 x=2-3(-6) = 2+18 = 20

OpenStudy (anonymous):